OpenStudy (anonymous):

Can anyone help with this 703(100)/h^2=20. Thanks

OpenStudy (campbell_st):

what do you need to do..?

OpenStudy (mertsj):

20h^2=703(100) h^2=703(5) h=+-59.3

OpenStudy (anonymous):

703(100)/20=h^2 h^2=3515 h=59.28

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Thank you.

OpenStudy (mertsj):

yw